OpenStudy (anonymous):

can you separate ln((x^2)/(x+3)) into lnx^2 + ln(x+3)?

OpenStudy (anonymous):

it would actually be 2lnx-ln(x+3)

OpenStudy (anonymous):

ln(x/y) = ln(x) - ln(y)

OpenStudy (anonymous):

okay! oh and another question. for integrals, you can separate it if it's an addition or subtraction between the numbers right?

OpenStudy (anonymous):

and ln(x^y) = yln(x)

OpenStudy (anonymous):

yeah, example \[\int\limits_{?}^{?}(x+5)dx = \int\limits_{?}^{?}xdx + \int\limits_{?}^{?}5dx\]

OpenStudy (anonymous):

thanks!